OpenStudy (anonymous):

Evaluate the expression: arcsin1/2

OpenStudy (anonymous):

arcsin(1/2)= π/6 radians

OpenStudy (anonymous):

Thanks. Can you show me how you got that?

OpenStudy (anonymous):

arcsin(1/2)= π/6 radians arcsin(1/2)= 5π/6 radians This is something that should be memorized. sin(π/6 )=1/2 If sin(π/6 )=1/2, arcsin(1/2) = π/6 radians sin(5π/6) is also 1/2 = This is 150 degrees so, arcsin(1/2) is also 5π/6

OpenStudy (anonymous):

Ahh okay thanks